# Invariants of Finite Orthogonal Groups of Plus Type in Odd Characteristic

### v1

@@ Abstract.
We describe the ring of invariants for the finite orthogonal group of plus type in odd characteristic acting on its defining representation. We also describe the invariants of its Sylow subgroup in the defining characteristic. In both cases we construct minimal algebra generating sets and describe the relations among the generators. Both rings of invariants are shown to be Complete Intersections and thus are Cohen-Macaulay. We expect the techniques we use will generalise to give a systematic computation for rings of invariants for all of the finite classical groups.
@@ 1. Introduction
There has been significant interest for a number of years in the rings of invariants of the defining representations of the finite classical groups. The rings of invariants for the symplectic groups were computed by David Carlisle and Peter Kropholler in the 1990s, see [1, §8.3]. The invariants in the unitary case were computed by Huah Chu and Shin-Yao Jow [9]. For the orthogonal groups, while there are computations in a number of special cases (see [7], [8], [10], [17] and [21] ) there is no general result.
We work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime. We compute the ring of invariants for the finite orthogonal groups ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ of plus type acting on its defining representation $V$ of dimension $n=2m$.
As an intermediate result, we compute the ring of invariants for a Sylow $p$-subgroup. For both of these rings we construct a minimal set of algebra generators and describe generators for the ideal of relations. Both of these rings are Complete Intersections, see Theorems 4.6 and 5.9, and are thus Cohen-Macaulay.
@@
We believe that the methods we use here generalise to give a systematic computation for rings of invariants for all of the finite classical groups; this is work in progress. For background on the finite classical groups we suggest [22]. For background on invariant theory see [1], [5], [11] or [20].
@@ 2. The Setting
For a symmetric bilinear form $\beta$ on a vector space $V$, the orthogonal group ${\rm O}(\beta)$ is the subgroup of the the general linear group ${\rm GL}(V)$ consisting of linear transformations $g$ satisfying $\beta(gv,gu)=\beta(v,u)$ for all $u,v\in V$. In this paper we work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime. In this context, if the dimension of $V$ is even, there are two equivalence classes of symmetric non-degenerate bilinear forms, one of plus type and one of minus type (see Chapter 11 of [22]). In this paper we focus on the plus type. We expect the generalisation of our methods to the minus type to be relatively straightforward.
@@
see [22, page 141]. For notational simplicity, we write $G_{m}:={{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. The ring of invariants is
@@ 3.7. New Invariants
For $v\in\mathcal{V}(x_{1},x_{2},\ldots,x_{m})$, choose $g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so that $gv\in V(x_{1},x_{2},\ldots,x_{m})$. From Theorem 3.9, $d_{i,m}\in S_{m}^{G_{m}}$. Therefore $d_{i,m}(gv)=d_{i,m}(v)$. Hence to show $\mathcal{V}(\mathcal{H})=\{\underline{0}\}$, it is sufficient to show $\mathcal{V}(x_{1},\ldots,x_{m},d_{1,m},\ldots,d_{m,m})=\{\underline{0}\}$. To do this we consider $d_{i,m}$ modulo the ideal $I:=\langle x_{1},x_{2},\ldots,x_{m}\rangle$.
@@ 4. Formulating the Main Theorem
Define $\nu:R_{n-1}\to\mathbb{Z}\cup\{\infty\}$ by $\nu(0)=\infty$ and, for $f\not=0$, $\nu(f)$ is the minimum value of ${R\mbox{-}\deg}$ on the non-zero terms of $f$. For example, $\nu(\xi_{1}^{q})=q$ and $\nu(\xi_{1}^{q}-\xi_{2}\xi_{1}^{(q-1)/2})=(q+1)/2$. Note that if $\nu(f)=\ell$, then $f\in(R_{n-1})_{+}^{\ell}\setminus(R_{n-1})_{+}^{\ell+1}$. It is easy to verify that for $f,h\in R_{n-1}$, we have $\nu(fh)=\nu(f)\nu(h)$ and $\nu(f+h)\geq{\rm min}\{\nu(f),\nu(h)\}$. It follows from [24, Ch IV, §9, Theorem 14], that defining $\nu(f/h)=\nu(f)-\nu(h)$ extends $\nu$ to a discrete valuation on $\mathcal{F}(R_{n-1})$. Note that the associated Discrete Valuation Ring, $\{a\in\mathcal{F}(R_{n-1})\mid\nu(a)\geq 0\}$, properly contains $R_{n-1}$, i.e., $R_{n-1}$ is not the associated Discrete Valuation Ring.
@@ 5. Sylow Invariants
The order of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is $2q^{m(m-1)}(q^{m}-1)\prod_{j=1}^{m-1}(q^{2j}-1)$, see [22, page 141]. Therefore a Sylow $p$-subgroup has order $q^{m(m-1)}$.
@@ 5.1. Computing the Hook Invariants
We work in the ring $\widetilde{T}:=T[a_{1},a_{2},\dots,a_{q}]$ where $a_{1},a_{2},\dots,a_{q}$ are $q$ new indeterminants. Write $f=\sum_{k}\sum_{j}c_{k,j}y_{1}^{k}t^{j}$. We have $\nu_{1}(c_{k,j})\geq||k||-1$ for all $k$ and $j$. Consider
@@
Then $\widetilde{f}$ is a linear combination of expressions of the form
@@
Terms involving $y_{1}^{k}u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}}$ occur multiple times in the expansion of $\widetilde{f}$ and collecting all these terms together we get an expression
$h_{k,i,j,\ell}(a_{1},a_{2},\dots,a_{q})y_{1}^{k}u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}}$
where $h_{k,i,j,\ell}\in\mathbb{F}_{q}[a_{1},a_{2},\dots,a_{q}]$ is homogeneous of degree $i+\ell$. From the definition of $\widetilde{f}$ we see that the polynomial $h_{k,i,j,\ell}(a_{1},a_{2},\dots,a_{q})$ must be invariant under any permutation of the $a_{s}$. Thus $h_{k,i,j,\ell}$ is a symmetric function. Now if we specialize the elements $a_{1},a_{2},\dots,a_{q}$ to be the $q$ elements in $\mathbb{F}_{q}$ then $\widetilde{f}$ specializes to $\mathcal{T}_{(y_{1},y,x,x_{1})}(f)$. Furthermore, under this specialization, $h_{k,i,j,\ell}(a_{1},a_{2},\dots,a_{q})$ must vanish unless its degree is a multiple of $q-1$ including the possibility that $h_{k,i,j,\ell}$ is constant. This condition on the degree of $h_{k,i,j,\ell}$ follows by noting that
$X^{q}-X=\prod_{a\in\mathbb{F}_{q}}(X-a)=\sum_{d=0}^{q}(-1)^{d}s_{d}(\mathbb{F}_{q})X^{q-d}$
where $s_{d}(\mathbb{F}_{q})$ is the $d^{\text{th}}$ elementary symmetric function in the $q$ elements of $\mathbb{F}_{q}$. In particular, $s_{d}(\mathbb{F}_{q})=0$ for all positive $d\neq q-1$.
The terms where $h_{k,i,j,\ell}$ has degree 0 correspond exactly to the terms in $f^{q}$. Since $||k^{q}||=||k||$ it is clear that $f^{q}$ is compliant and thus the terms arising from $h_{k,i,j,\ell}$ being constant are not an obstruction to $\widetilde{f}$ being compliant. Thus we may restrict our attention to terms where $\deg(h_{k,i,j,\ell})$ is a positive multiple of $q-1$. But if $i+\ell$ is a positive multiple of $q-1$ then so is $||i+\ell||$ by Lemma 5.6. In particular $||i+\ell||\geq q-1$
@@
Thus $\nu_{1}(u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}})\geq\ell+\sum_{s=1}^{q}(||k_{s}||-1)\geq||k||-1$ as required. ∎
@@ Acknowledgement
The computer algebra package [2] was very useful in this work. It was used to test and confirm many hypotheses in low dimensions for small primes. We are grateful to Karl Dilcher and Keith Taylor (Canadian Mathematical Society book editors) and Donna Chernyk (Springer) for encouraging us to work on this problem.

### v2

@@ Abstract.
We describe the rings of invariants for the finite orthogonal groups of plus type in odd characteristic acting on the defining representations. We also describe the invariants of the corresponding Sylow subgroups in the defining characteristic. In both cases we construct minimal algebra generating sets and describe the relations among the generators. Both rings of invariants are shown to be complete intersections and thus are Cohen-Macaulay. We expect the techniques we use will generalise to give a systematic computation for rings of invariants for all of the finite classical groups in odd characteristic.
@@ 1. Introduction
The fundamental problem in the invariant theory of finite groups is to determine the ring of invariants of a representation of a finite group. The celebrated theorem of Shepherd–Todd and Chevalley proves that such a ring of invariants in characteristic zero is a polynomial ring if and only if the group action is generated by pseudo-reflections (group elements which fix pointwise a hyperplane). An excellent survey of the characteristic zero results and theory in characteristic zero is the article by Stanley ([22]).
In positive characteristic the situation is much more complex. In 1911, L.E. Dickson [12] gave an explicit description of the ring of invariants of the general linear group over any finite field. In the past few decades there has been significant progress in determining the rings of invariants of the defining representations of the finite classical groups. The rings of invariants for the symplectic groups were computed by David Carlisle and Peter Kropholler in the 1990s, see [1, §8.3]. The invariants in the unitary case were computed by Huah Chu and Shin-Yao Jow [9]. For the orthogonal groups, while there are computations in a number of special cases (see [7], [8], [10], [17] and [21] ) there is no general result. While almost all of the defining representations for these groups are generated by pseudo-reflections, the rings of invariants are rarely polynomial rings.
We work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime. We compute the ring of invariants for the finite orthogonal group ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ of plus type acting on its defining representation $V$ of dimension $n=2m$.
As an intermediate result, we compute the ring of invariants for a Sylow $p$-subgroup. For both of these rings we construct a minimal set of algebra generators and describe generators for the ideal of relations. Both of these rings are complete intersections, see Theorems 4.6 and 5.9, and are thus Cohen-Macaulay. We note that for modular representations, rings of invariants are rarely Cohen-Macaulay, let alone complete intersections.
@@
We believe that the methods we use here generalise to give a systematic computation for rings of invariants for all of the finite classical groups, at least for odd characteristic; this is work in progress. For background on the finite classical groups we suggest [23]. For background on invariant theory see [1], [5], [11] or [20].
@@ 2. The Setting
For a symmetric bilinear form $\beta$ on a vector space $V$, the orthogonal group ${\rm O}(\beta)$ is the subgroup of the the general linear group ${\rm GL}(V)$ consisting of linear transformations $g$ satisfying $\beta(gv,gu)=\beta(v,u)$ for all $u,v\in V$. In this paper we work over the finite field $\mathbb{F}_{q}$ of order $q$ and characteristic $p$ where $p$ is an odd prime.
In this context, if the dimension of $V$ is even, there are two equivalence classes of symmetric non-degenerate bilinear forms, one of plus type and one of minus type (see Chapter 11 of [23]). In this paper we focus on the plus type. We expect the generalisation of our methods to the minus type to be relatively straightforward.
@@
see [23, page 141]. For notational simplicity, we write $G_{m}:={{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$. The ring of invariants is
@@ 3.7. New Invariants
For $v\in\mathcal{V}(\xi_{0},\ldots,\xi_{m-1})$, choose $g\in{{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ so that $gv\in\mathcal{V}(x_{1},x_{2},\ldots,x_{m})$. From Theorem 3.9, $d_{i,m}\in S_{m}^{G_{m}}$. Therefore $d_{i,m}(gv)=d_{i,m}(v)$. Hence to show $\mathcal{V}(\mathcal{H})=\{\underline{0}\}$, it is sufficient to show $\mathcal{V}(x_{1},\ldots,x_{m},d_{1,m},\ldots,d_{m,m})=\{\underline{0}\}$. To do this we consider $d_{i,m}$ modulo the ideal $I:=\langle x_{1},x_{2},\ldots,x_{m}\rangle$.
@@ 4. Formulating the Main Theorem
Define $\nu:R_{n-1}\to\mathbb{Z}\cup\{\infty\}$ by $\nu(0)=\infty$ and, for $f\not=0$, $\nu(f)$ is the minimum value of ${R\mbox{-}\deg}$ on the non-zero terms of $f$. For example, $\nu(\xi_{1}^{q})=q$ and $\nu(\xi_{1}^{q}-\xi_{2}\xi_{1}^{(q-1)/2})=(q+1)/2$. Note that if $\nu(f)=\ell$, then $f\in(R_{n-1})_{+}^{\ell}\setminus(R_{n-1})_{+}^{\ell+1}$. It is easy to verify that for $f,h\in R_{n-1}$, we have $\nu(fh)=\nu(f)\nu(h)$ and $\nu(f+h)\geq{\rm min}\{\nu(f),\nu(h)\}$. It follows from [25, Ch IV, §9, Theorem 14], that defining $\nu(f/h)=\nu(f)-\nu(h)$ extends $\nu$ to a discrete valuation on $\mathcal{F}(R_{n-1})$. Note that the associated Discrete Valuation Ring, $\{a\in\mathcal{F}(R_{n-1})\mid\nu(a)\geq 0\}$, properly contains $R_{n-1}$, i.e., $R_{n-1}$ is not the associated Discrete Valuation Ring.
@@ 5. Sylow Invariants
The order of ${{\text{O}^{+}_{{2m}}(\mathbb{F}_{q})}}$ is $2q^{m(m-1)}(q^{m}-1)\prod_{j=1}^{m-1}(q^{2j}-1)$, see [23, page 141]. Therefore a Sylow $p$-subgroup has order $q^{m(m-1)}$.
@@ 5.1. Computing the Hook Invariants
We work in the ring $\widetilde{T}:=T[a_{1},a_{2},\dots,a_{q}]$ where $a_{1},a_{2},\dots,a_{q}$ are $q$ new indeterminants. Write $f=\sum_{k}\sum_{j}c_{k,j}y_{1}^{k}t^{j}$. We have $\nu(c_{k,j})\geq||k||-1$ for all $k$ and $j$. Consider
@@
Thus $\widetilde{f}$ is a linear combination of expressions of the form
@@
Clearly $\widetilde{f}$ is invariant under any permutation of the $a_{s}$. Thus we may express $\widetilde{f}$ using the elementary symmetric functions $E_{i}(a_{1},a_{2},\dots,a_{q})$ of degree $i$ in the $a_{i}$, i.e., $\widetilde{f}=\sum_{\vec{e}}h_{(e_{1},e_{2},\dots,e_{q})}E_{1}^{e_{1}}E_{2}^{e_{2}}\cdots E_{q}^{e_{q}}$ with $h_{(e_{1},e_{2},\dots,e_{q})}\in T$. When we specialize the $a_{i}$ to the elements of the field $\mathbb{F}_{q}=\{\alpha_{1},\alpha_{2},\dots,\alpha_{q}\}$, $\widetilde{f}$ specializes to $\mathcal{T}_{(y_{1},u,t,x_{1})}(f)$.
Note that $E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})=0$ for $i\neq q-1$ and $i\geq 1$. This follows from
$X^{q}-X=\prod_{a\in\mathbb{F}_{q}}(X-a)=\sum_{i=0}^{q}(-1)^{i}E_{i}(\alpha_{1},\alpha_{2},\dots,\alpha_{q})X^{q-i}$
Therefore in the expression for $\widetilde{f}$ it suffices to consider only terms of the form $h_{(0,0,\dots,0,e,0)}E_{q-1}^{e}$.
Terms occurring in $h_{(0,0,\dots,0,e,0)}$ satisfy $i+\ell=e(q-1)$. The term $h_{(0,0,\dots,0,0,0)}$ when $i=\ell=0$ is just $h_{(0,0,\dots,0,0,0)}=\prod_{s=1}^{q}f(y_{1},t)=f(y_{1},t)^{q}$. Since $||k^{q}||=||k||$ it is clear that $h_{(0,0,\dots,0,0,0)}$ is compliant.
@@
For $h_{(0,0,\dots,0,e,0)}$ with $e\geq 1$ we have $i+\ell=e(q-1)$ is a positive multiple of $q-1$. Thus $||i+\ell||$ is also a positive multiple of $q-1$ by Lemma 5.6. In particular $||i+\ell||\geq q-1$.
@@
Thus $\nu(u^{i}t^{j}x_{1}^{\ell}\prod_{s=1}^{q}c_{k_{s},j_{s}})\geq\ell+\sum_{s=1}^{q}(||k_{s}||-1)\geq||k||-1$ as required. ∎
@@ Acknowledgement
The computer algebra package Magma [2] was very useful in this work. It was used to test and confirm many hypotheses in low dimensions for small primes. We are grateful to Karl Dilcher and Keith Taylor (Canadian Mathematical Society book editors) and Donna Chernyk (Springer) for encouraging us to work on this problem.